The ages, in years of three sisters are consecutive odd integers. Three years ago, the oldest sister was three times as old as the youngest. How old is each now
How do we set this problem up: We've tried several equations, but we're missing something. Please help. Both my husband and I have tried to help our daughter, We were given the answers of 5, 7, 9. Please help us set up the equation and show steps to solving. Thanks so much

Answer by scott8148(5880)

consecutive odd integers are separated by 2; so the ages are x, x+2, x+4 ... three years ago is -3

(x+4)-3=3((x)-3) ... x+1=3x-9 ... 10=2x ... 5=x ... the ages are 5, 7, 9